# Overcoming label shift in targeted federated learning

### v1

@@ 6 Discussion
We explored targeted federated learning under label shift, a scenario where client data distributions differ from a target domain with a known label distribution, but no target samples are available. We demonstrated that traditional approaches, such as FedAvg, which assume identical distributions between clients and the target, fail to adapt effectively in this context due to biased aggregation of client updates. To address this, we proposed FedPALS, a novel aggregation strategy that optimally combines client updates to align with the target distribution, ensuring that the aggregated model minimizes target risk. Empirically, across diverse tasks, we showed that under label shift, FedPALS significantly outperforms standard methods like FedAvg, FedProx and SCAFFOLD. Specifically, when the target label distribution lies within the convex hull of the client distributions, FedPALS finds the solution with the largest effective sample size, leading to a model that is most faithful to the target distribution. More generally, FedPALS balances the trade-off between matching the target label distribution and minimizing variance in the model updates. Our experiments further highlight that FedPALS excels in challenging scenarios where label sparsity and client heterogeneity hinder the performance of conventional federated learning methods.
We also observed that the choice of the trade-off parameter $\lambda$ is crucial for achieving optimal performance in tasks such as iWildCam, where the label shift assumption may not fully hold. Interestingly, the early performance gains observed during training suggest that dynamically tuning $\lambda$ over time could enhance performance of FedPALS. A promising avenue for future work would be exploring adaptive strategies for dynamically tuning $\lambda$. Additionally, our weighting approach could be extended to the covariate shift setting, where input distributions vary between clients and the target. This extension is feasible when the central server has access to unlabeled target and client samples, akin to unsupervised domain adaptation in the centralized learning paradigm.
Acknowledgements
AB and FDJ were supported by the Wallenberg AI, Autonomous Systems and Software Program (WASP), funded by the Knut and Alice Wallenberg Foundation.
The computations and data handling were enabled by resources provided by the National Academic Infrastructure for Supercomputing in Sweden (NAISS), partially funded by the Swedish Research Council through grant agreement no. 2022-06725.
We also thank Samuel Håkansson, Anton Matsson, Lena Stempfle, Herman Bergström, Javad Aliakbari, Olof Mogren, John Martinsson, Aleksis Pirinen and Martin Willbo for providing feedback on earlier drafts of this work.
Appendix
@@ Appendix A Experimental details
Here we provide additional details about the experimental setup for the different tasks.
A.1 Oracle construction
@@
A.2 Synthetic task
@@ A.3 PACS
We pick four values of $\lambda$, [0,12,42,93], which approximately correspond to an ESS of $15\%,25\%,50\%$ and $75\%$ respectively. We use the same hyperparameters during training as Bai et al. [2024] report using in their paper.
A.4 iWildCam
We perform this experiment using the methodology described in Bai et al. [2024] with the heterogeneity set to the maximum setting, i.e., $\lambda=0$ in their construction. [^3] Note that this is not the same $\lambda$ used in the trade-off in FedPALS. We use the same hyperparameters which is used for FedAvg in the same work to train FedPALS. We perform 80 rounds of training and, we then select the best performing model based on held out validation performance and report the mean and standard deviation over three random seeds. This can be seen in Table 2. We pick four values of $\lambda$, [0,600,2500,5800], which approximately correspond to an ESS of $8\%,25\%,50\%$ and $75\%$ respectively.
@@ C.2 FedPALS in the limits
[…]
We have the additional conditions of primal feasibility, i.e.
@@
Using complementary slackness, $\tau g(\alpha^{*})=0$, and that $\tau\geq 0$ we find that $\tau=0$. Thus we can write (15) as
$2\sum_{i}\frac{\alpha_{i}}{n_{i}}-\mu=0\implies\mu=2\sum_{i}\frac{\alpha^{*}_{i}}{n_{i}}$
Combining this with the primal feasibility condition $h(\alpha^{*})=0$ we obtain $\mu=\frac{2}{\sum_{i}n_{i}}$ which implies $\alpha_{i}^{*}=\frac{n_{i}}{\sum_{i}n_{i}}$. This proves also Proposition 2. ∎

### v2

@@ 6 Discussion
We have explored target-aware federated learning under label shift, a scenario where client data distributions differ from a target domain with a known label distribution, but no target samples are available. We demonstrated that traditional approaches, such as federated averaging (FedAvg), which assume identical distributions between the client aggregate and the target, fail to adapt effectively in this context due to biased aggregation of client updates. To address this, we proposed FedPALS, a novel aggregation strategy that optimally combines client updates to align with the target distribution, ensuring that the aggregated model minimizes target risk. Empirically, across diverse tasks, we showed that under label shift, FedPALS significantly outperforms standard methods like FedAvg, FedProx, FedRS and SCAFFOLD, as well as AFL. Specifically, when the target label distribution lies within the convex hull of the client distributions, FedPALS finds the solution with the largest effective sample size, leading to a model that is most faithful to the target distribution. More generally, FedPALS balances the trade-off between matching the target label distribution and minimizing variance in the model updates. Our experiments further highlight that FedPALS excels in challenging scenarios where label sparsity and client heterogeneity hinder the performance of conventional federated learning methods.
One of the limitations of FedPALS is label shift–the assumption that label-conditional input distributions are equal in all clients and the target. We observed empirically that selecting the trade-off parameter $\lambda$ is crucial for optimal performance in tasks such as iWildCam, where this assumption may not fully hold. Future work should aim to detect and overcome violations of this assumption.
@@ Appendix A Experimental details
Here we provide additional details about the experimental setup for the different tasks. The code will be made available upon acceptance.
A.1 Choice of hyperparameter $\lambda$
A salient question in Cases 2 & 3, as discussed in Section 3.1, is how to choose the strength of the regularization, $\lambda$. A larger value will generally favor influence from more clients, provided that they have sufficiently many samples. When $T\not\in\mathrm{Conv}(S)$, the convex combination closest to $T$ could have weight on a single vertex. This will likely hurt the generalizability of the resulting classifier. In experiments, we compare values of $\lambda$ that yield different effective sample sizes, such as 10%, 25%, 50% or 75% of the original sample size, $N$. We can find these using binary search by solving (9) and calculate the ESS. One could select $\lambda$ heuristically based on the the ESS, or treat $\lambda$ as a hyperparameter and select it using a validation set. Although this would entail training and evaluating several models which can be seen as a limitation. We elect to choose a small set of $\lambda$ values based on the ESS heuristic and train models for these. Then we use a validation set to select the best performing model. This highlights the usefulness of the ESS as a heuristic. If it is unclear which values to pick, one could elect for a simple strategy of taking the ESS of $\lambda=0$ and 100% and taking $l$ equidistributed values in between the two extremes, for some small integer $l$.
A.2 Sampling strategies
Sparsity sampling entails randomly removing a subset of labels from clients and the target domain following the data set partitioning technique introduced in McMahan et al. [2017]. Each client is assigned $C$ random labels, with an equal number of samples for each label and no overlap among clients. Sparsity sampling has been extensively used in subsequent studies [Geyer et al., 2017, Li et al., 2020a, 2022].
Dirichlet sampling simulates non-i.i.d. label distributions by, for each client $i$, drawing a sample $p_{i}\sim\mathrm{Dirichlet}_{K}(\beta)$, where $p_{i}(k)$ represents the proportion of samples in client $i$ that have label $k\in[K]$. The concentration parameter $\beta>0$ controls the sparsity of the client distributions. In dirichlet sampling, using a smaller $\beta$ results in more heterogeneous client data sets, while a larger value approximates an i.i.d. setting. This widely-used method for sampling clients was first introduced by Yurochkin et al. [2019].
A.3 Oracle construction
@@
A.4 Perturbation of target marginal $\mathcal{T}$
In an experiment we perturb the given target label marginals, $\mathcal{T}$, to evaluate the performance impact of noise in the estimate. We do this by generating gaussian noise, $\epsilon$, and then we add the noise to the label marginal to create a new target $\mathcal{T}_{p}$. We modulate the size of the noise with a parameter $\delta$ and only add the positive noise values.
$\mathcal{T}_{p}=\mathcal{T}+\delta\max(\epsilon,0)$
This is then normalised and used as the new target label marginal. This perturbation was done on the PACS experiment with $\delta\in[10^{-3},10^{-2},5\times 10^{-1}]$ and repeated for three seeds. The results are given in Table 2 where we see that the performance decreases with increasing noise.
$\delta$ Accuracy
$10^{-3}$ 88.8
$10^{-2}$ 85.2
$5\times 10^{-1}$ 81.4
Table 2: Results of perturbing $\mathcal{T}$ with varying noise levels $\delta$.
A.5 Synthetic task
@@ A.6 PACS
We pick four values of $\lambda$, [0,12,42,93], which approximately correspond to an ESS of $15\%,25\%,50\%$ and $75\%$ respectively. We use the same hyperparameters during training as Bai et al. [2024] report using in their paper. Furthermore, we use the cross entropy loss in this task.
A.7 iWildCam
We perform this experiment using the methodology described in Bai et al. [2024] with the heterogeneity set to the maximum setting, i.e., $\lambda=0$ in their construction. [^1] Note that this is not the same $\lambda$ used in the trade-off in FedPALS. We use the same hyperparameters which is used for FedAvg in the same work to train FedPALS. We perform 80 rounds of training and, we then select the best performing model based on held out validation performance and report the mean and standard deviation over three random seeds. This can be seen in Table 3. We pick four values of $\lambda$, [0,600,2500,5800], which approximately correspond to an ESS of $8\%,25\%,50\%$ and $75\%$ respectively. We use the cross entropy loss in this task.
@@ B.3 Local epochs and number of clients
B.4 Synthetic experiment: effect of projection distance on test error
@@ C.2 FedPALS in the limits
$\nabla_{\alpha}\mathcal{L}(\alpha^{*})=0\implies\forall i:\leavevmode\nobreak\ \leavevmode\nobreak\ 2\frac{\alpha^{*}_{i}}{n_{i}}-\mu-\tau=0\leavevmode\nobreak\ .$ (17)
In other words, the following ratio is a constant,
$\forall i\quad\frac{\alpha^{*}_{i}}{n_{i}}=c$
for some constant $c$. We have the additional conditions of primal feasibility, i.e.
@@
From the first constraint, we have $\sum_{i=1}^{M}\alpha^{*}_{i}=1$, and thus,
$\sum_{i=1}^{M}\alpha_{i}^{*}=c\sum_{i=1}^{M}n_{i}=1$
which implies that $c=1/\sum_{i=1}^{M}n_{i}$ and thus
$\forall i:\alpha_{i}^{*}=\frac{n_{i}}{\sum_{i=1}^{M}n_{i}}\leavevmode\nobreak\ .$